OpenStudy (anonymous):

How to solve simultaneously: x^4-6x^2+60x+36=68x+39

OpenStudy (paki):

combine the like terms ans solve...

OpenStudy (anonymous):

yeah and?

OpenStudy (anonymous):

i did combine both and equaled the equation to zero

OpenStudy (paki):

that will be the answer....

OpenStudy (anonymous):

i need to factorise the equation

OpenStudy (anonymous):

how to factorise a quartic

OpenStudy (paki):

yup i know... then use a quadratic formula and solve...

OpenStudy (anonymous):

the x-coordinate of the point it intersect is -1 so the when factorising it: (x+1)... and i don't know the rest

OpenStudy (paki):

pk tell me, how you simplified this for factors....?

OpenStudy (anonymous):

do you know how to factorise the rest?

OpenStudy (anonymous):

\[x^4-6x^2-8x-3=0,put x=-1,1-6+8-3=0,0=0,so x+1 is a factor\] divide synthetically by -1 or divide by x+1 long division.

OpenStudy (anonymous):



OpenStudy (tkhunny):

What does "simultaneously" mean with only one equation?